OpenStudy (joshoyen):

Solve by isolating the x^2 term and taking the square roots ( + and - ). 1. x^2 = 25 2. x^2 - 16 = 0 3. -2x^2 + 40 = 0

OpenStudy (danjs):

recall \[\sqrt{x^2}=\left| x \right|\]

OpenStudy (joshoyen):

oh so you just find the square root of 25

OpenStudy (joshoyen):

?

OpenStudy (joshoyen):

how about x^2 - 16 = 0, do you add 16 to both sides?

OpenStudy (danjs):

yes, square root both sides of the =, and remember the abs value

OpenStudy (danjs):

abs value of x = 5, so x can be + or - 5

OpenStudy (joshoyen):

oh so its x = |5|

OpenStudy (danjs):

same for the second, isolate the x^2, take the square root of everything

OpenStudy (joshoyen):

square root of 16 is 4

OpenStudy (danjs):

yes, and the square root of x^2 = absolute value of x because x^2 will be the same for the + or - value of x

OpenStudy (joshoyen):

ahh i see alright thanks!, now how abbout the 3rd one? what step do i do first

OpenStudy (danjs):

\[\sqrt{x^2}=\sqrt{16}\] \[\left| x \right|=4\] x=+ 4 or x=-4

OpenStudy (joshoyen):

gotcha

OpenStudy (danjs):

try the last one

OpenStudy (joshoyen):

is it x^2 - 20 = 0

OpenStudy (danjs):

ok, so you divided everything by -2

OpenStudy (joshoyen):

yep

OpenStudy (danjs):

k, same as #2 prob now

OpenStudy (joshoyen):

ohh okay i see now thanks!

OpenStudy (danjs):

you can do whatever operation you want, as long as it is to both sides of the equation

OpenStudy (joshoyen):

alright gotcha tyvm

OpenStudy (danjs):

i just think move the 40 over by subtracting 40 isolate x^2 by dividing everything by -2 square root everything

OpenStudy (danjs):

welcome